OpenStudy (anonymous):

solve the equaion: 3x+2div x-2+ 1div x= -2div x ^{2}-2x

OpenStudy (anonymous):

\[3x+2\div x-2+ 1\div x= -2\div x ^{2}-2x\] considering there are brackets (I don't see any), \[(3x+2)/(x-2) + 1/x = -2/(x ^{2}-2x)\] reduce to a common denominator \[(3x ^{2}+3x-2)/(x ^{2}-2x) = -2/(x ^{2}-2x)\] multiply their denominator to each side and \[3x ^{2}+3x-2=-2\] is what you get. add 2 to each side \[3x ^{2}+3x=0\] divide by 3, and \[x(x+1)=0\] is what is left. \[\therefore x=0,(-1)\]

jhonyy9 (jhonyy9):

3x+2 1 -2 ---- + ---- = --------- x-2 x x*(2) - 2x 3x+2 1 -2 ----- + ---- = ------- x-2 x x(x-2) x(3x+2) + x-2 = -2 3x*2 +2x +x -2= -2 3x*2+3x=0 3x(x+1)=0 3x=0 --- x=0 x+1=0 -- x= -1

OpenStudy (nikvist):

x=0 is not solution of this equation